OpenStudy (mollylasko):

Crop researchers plant 15 plots with a new variety of corn. The yields in bushels per acre are: 138.0 139.1 113.0 132.5 140.7 109.7 118.9 134.8 109.6 127.3 115.6 130.4 130.2 111.7 105.5 Assume that Math Notation bushels per acre. a) Find the 90% confidence interval for the mean yield for this variety of corn. b) Find the 95% confidence interval. c) Find the 99% confidence interval. d) How do the margins of error in (a), (b), and (c) change as the confidence level increases?

OpenStudy (mollylasko):

@ParthKohli

OpenStudy (mollylasko):

@jim_thompson5910

OpenStudy (mollylasko):

@jigglypuff314

jimthompson5910 (jim_thompson5910):

Are you familiar with the confidence interval formula?

OpenStudy (mollylasko):

no @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so you've never seen these formulas before? or anything like it? \[\Large L = \bar{x} - z^{*} \frac{s}{\sqrt{n}}\] \[\Large U = \bar{x} + z^{*} \frac{s}{\sqrt{n}}\]

OpenStudy (mollylasko):

no :/ flvs kinda sucks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok let's go over the formulas then. They aren't too bad the first symbol that we'll go over is this \[\LARGE \bar{x}\] that's pronounced "x-bar" and it is the sample mean to find the sample mean, you will step 1) add up the values step 2) divide the result (from step 1) by the number of values tell me what you get

OpenStudy (mollylasko):

123.8 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

very good. Ok onto the next symbol \(\LARGE z^{*}\) which is often pronounced "z-star". This is the critical z-value based on the confidence level It is impossible to calculate this by hand (well it is possible but it will take a very long time). So we can use a calculator or a table Let's use the table https://www.ltcconline.net/greenl/courses/201/estimation/smallConfLevelTable.htm when the confidence level is 0.90, what is the value of z?

OpenStudy (mollylasko):

1.645? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes, so \(\LARGE z^{*} = 1.645\) for the 90% confidence level

jimthompson5910 (jim_thompson5910):

the 's' in the formula refers to the sample standard deviation. Let's use a calculator to compute this if you don't have a calc that can handle standard deviation, then use this one http://www.miniwebtool.com/sample-standard-deviation-calculator/

OpenStudy (mollylasko):

so is that my a? the 1.645? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no, we're still working on the pieces of the formula

jimthompson5910 (jim_thompson5910):

part a) isn't done yet

jimthompson5910 (jim_thompson5910):

here's how you can enter the data (see attached)

OpenStudy (mollylasko):

and for the sample standard deviation i got 12.2603891805 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

ok so s = 12.260389 approximately

jimthompson5910 (jim_thompson5910):

finally the n is simply the sample size. No need to use a table or calc here. We just look at the problem it says `Crop researchers plant 15 plots ` so n = 15

jimthompson5910 (jim_thompson5910):

let's plug these values in to get... \[\Large L = \bar{x} - z^{*} \frac{s}{\sqrt{n}}\] \[\Large L = 123.8 - 1.645* \frac{12.260389}{\sqrt{15}}\] \[\Large L = ???\] Use a calculator to compute the expression and tell me what you get

OpenStudy (mollylasko):

i received 386.6961679 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's incorrect. How did you enter that into your calc?

OpenStudy (mollylasko):

just how you wrote it, but let me try again @jim_thompson5910

OpenStudy (mollylasko):

i got 118.592557 this time idk what happened the last time lol @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you can try this calc http://web2.0calc.com/ and type in `123.8 - 1.645*(12.260389/sqrt(15))`

jimthompson5910 (jim_thompson5910):

much better, yes L = 118.592557 approx

jimthompson5910 (jim_thompson5910):

this is the lower bound of the confidence level L = Lower bound U is the upper bound so let's calculate that now \[\Large U = \bar{x} + z^{*} \frac{s}{\sqrt{n}}\] \[\Large U = 123.8 + 1.645* \frac{12.260389}{\sqrt{15}}\] \[\Large U = ???\]

OpenStudy (mollylasko):

is the upper one 129.007443? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

so the 90% confidence interval stretches from 118.592557 to 129.007443

OpenStudy (mollylasko):

is that a? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

let's round each value to 1 decimal place to get 118.6 and 129.0 so the 90% confidence interval is (118.6, 129.0) agreed?

jimthompson5910 (jim_thompson5910):

I'm rounding to 1 decimal place because the other data values are given to 1 decimal place. Your book or teacher may want you to round to 2 decimal places

OpenStudy (mollylasko):

okay :)

jimthompson5910 (jim_thompson5910):

yes once you get to (118.6, 129.0) or (118.59, 129.01), then you're done with part a)

OpenStudy (mollylasko):

yay!! now b lol @jim_thompson5910

jimthompson5910 (jim_thompson5910):

part b) is going to be the same as part a) but the only difference is that z value will change use the table and tell me what z is equal to when the confidence level is 95%

OpenStudy (mollylasko):

1.96 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

repeat part a) but just use z = 1.96 instead

jimthompson5910 (jim_thompson5910):

xbar, s and n will remain the same

OpenStudy (mollylasko):

for te U i got 130.0046129 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good how about L?

OpenStudy (mollylasko):

117.5953871 ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

if we round to 1 decimal place, the CI (confidence interval) is (117.6, 130.0) if we round to 2 decimal places, the CI is (117.60, 130.00) so effectively the two results are identical

OpenStudy (mollylasko):

yayy @jim_thompson5910

jimthompson5910 (jim_thompson5910):

for part c) you'll do the same as a) & b) but as you can probably guess by now, the z value will change. xbar, s, & n will stay the same

OpenStudy (mollylasko):

sweet this is so much easier than i thought it would be thank you !! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

then in part d), you just make an observation of what is happening (as the confidence level increases, do the intervals increase in size? or do the intervals decrease in size? or do the intervals stay the same?)

jimthompson5910 (jim_thompson5910):

or specifically you focus on the margin of error